dy/dx-y= e^x y^2

1
Expert's answer
2022-06-06T14:34:03-0400
"y'-y=e^xy^2"




"\dfrac{y'}{y^2}-\dfrac{1}{y}=e^x"

Use the change of variable

"z=y^{1-2}=y^{-1}"

"z=y^{-1}"

Differentiate both sides with respect to "x"


"z'=-y^{-2}y'"

Then


"y'=-z'y^2=-z'z^{-2}"

Substitute


"z'+z=-e^x"

We get the linear equation for the function "z(x)." To solve it, we use the integrating factor:


"u(x)=e^{\int(1)dx}=e^x"

"e^xz'+e^xz=-e^{2x}"

"d(ze^x)=-e^{2x}dx"

Integrate


"\int d(ze^x)=-\int e^{2x}dx"

"ze^x=-\dfrac{1}{2}e^{2x}+\dfrac{1}{2}C"

"z=-\dfrac{1}{2}e^x+\dfrac{1}{2}Ce^{-x}"

Then


"y=\dfrac{1}{-\dfrac{1}{2}e^x+\dfrac{1}{2}Ce^{-x}}, y\not=0"

"y=\dfrac{2}{-e^x+Ce^{-x}}, or\ y=0"